OpenStudy (anonymous):

medal & fan :) Find the length of the unknown side. Round your answer to the nearest tenth.

OpenStudy (anonymous):



OpenStudy (anonymous):

a^2 + b^2 = c^2 18^2 + b^2 = 75^2 36 + b^2 = 150 150-36 = 114 Your answer is 114

OpenStudy (anonymous):

@KngTHnd3r it can't be bc, its not in my answer choices :p 65.4 70.5 72.8 93.0

OpenStudy (anonymous):

What Unit is this?

hartnn (hartnn):

\(\Large Pythagoras~ theorem \\\Large x^2+18^2=75^2 \\\Large x^2+324 =5625 \)

hartnn (hartnn):

find x